OpenStudy (dls):

Simplify this! \large \tan^{-1} \frac{ \sqrt{1+x^2}-1}{x},x \neq 0

OpenStudy (dls):

\[\large \tan^{-1} \frac{ \sqrt{1+x^2}-1}{x},x \neq 0\]

OpenStudy (dls):

Master @ParthKohli

Parth (parthkohli):

Thanks for calling me, but I suck at trig :-P

OpenStudy (dls):

Hmm,nevermind tag someone better <3

Parth (parthkohli):

I have called the real master here.

OpenStudy (dls):

I await to see him.

Parth (parthkohli):

He has seen my FB message. Let's see when he's here.

Parth (parthkohli):

He says he's coming.

OpenStudy (dls):

I await I await!

OpenStudy (dls):

Lets see who it is D:

OpenStudy (dls):

oooooo :")

Parth (parthkohli):

Told you.

mathslover (mathslover):

Got to go , I am giving you a hint : \[\tan ^{-1 } (\cfrac{\sqrt{1+x^2} - 1}{\sqrt{1+x^2-1}})\]

OpenStudy (dls):

\[\LARGE \tan^{-1} \frac{\sec \theta-1}{\tan \theta}=> \frac{1-\cos \theta}{\sin \theta}\] used half angle formula and got it :

OpenStudy (dls):

=)

mathslover (mathslover):

Will this property help you?

OpenStudy (dls):

@mathslover I just substituted x=tan theta

mathslover (mathslover):

Ok!

OpenStudy (anonymous):

multiply both numerator and denominator by root(1+x^2) +1

OpenStudy (anonymous):

Then use the property given by @mathslover

OpenStudy (dls):

yo!

Parth (parthkohli):

I don't fit in this genius world.

OpenStudy (dls):

but substitution method is better :P

OpenStudy (anonymous):

@ParthKohli may be because u r better than genius

OpenStudy (dls):

^^

mathslover (mathslover):

:P if you love substitution or you think it is better or simpler then go for it but do try the another method also. It would be beneficial for you.

Parth (parthkohli):

I am overrated -_-